ROTATIONAL SPECTRA OF JET-COOLED ION COMPLEXES

\underline{YASUHIRO OHSHIMA}, Department of Chemistry, Graduate School of Science, Kyoto University, Sakyo-ku, Kyoto 606-01, JAPAN; YOSHIHIRO SUMIYOSHI, AND YASUKI ENDO, Department of Pure and Applied Sciences, Faculty of Arts and Sciences, The University of Tokyo, Meguro-ku, Tokyo 153, JAPAN.

Microwave spectra of complexes bearing molecular ions, Ar-D_{3}^{+}/HD_{2}^{+} and Ar-HCO^{+}, were explored in a jetcooled condition. The ionic complexes were formed by a pulsed electric discharge in a supersonic expansion. Their transitions were observed in the centimeter-wave region (7-40 GHz) with a Fourier-transform microwave spectrometer, guided by the previous millimeter-wave study in a static cell^{a} or the recent infrared photodissociation study in a mass-selected ion beam.^{b} For Ar-HCO^{+}, rotational transitions of mono-substituted species on H, C, and O were also measured. The isotopic substitution has confirmed the linear Ar⋯H-C-O configuration with the Ar⋯H distance of 2.13 Å. The centrifugal-distortion constant for Ar-HCO^{+} is two order of magnitude smaller than that for Ar-HCN, indicating much stronger binding in the ionic complex than the isoelectronic neutral counterpart.

aM. Bogey, H. Bolvin, C. Demuynck, and J. L. Destombes, Phys. Rev. Lett. \underline{58}, 988 (1987).

bS. A. Nizkorodov, O. Dopfer, T. Ruchti, M. Meuwly, J. P. Maier, and E. J. Bieske, J. Phys. Chem. \underline{99}, 17118 (1995).